OpenStudy (anonymous):

Consider two parallel line segments. Part A: Find the slope of segment CD with endpoints C(−7, −2) and D(−4, 2) Answer is 4/3 Part B: What is the value of y so that segment AB with endpoints A(4, y) and B(7, 4) is parallel to segment CD?

OpenStudy (anonymous):

@mehek14

Mehek (mehek14):

Part A is correct

OpenStudy (anonymous):

What about Part b? @Mehek14

Mehek (mehek14):

part b \(\dfrac{4-y}{7-4}=\dfrac{4}{3}\)

Mehek (mehek14):

4 - y = 4 y = ?

OpenStudy (anonymous):

y = 0?

Mehek (mehek14):

correct

OpenStudy (anonymous):

Thank you!

Mehek (mehek14):

ur welcome